# An Adaptive Algorithm for Bilevel Optimization on Riemannian Manifolds

### v2

@@ G.2.1 Improvement on objective function for one step update
where (a) follows from Baillon-Haddad Theorem [3, Theorem 5.8 (iv)], i.e., for $y\in{\mathcal{M}}_{y}$ and $v\in T_{y}{\mathcal{M}}_{y}$, it holds that
@@
where (a) follows from (32), (b) follows from $c_{n}>C_{c}\geq L_{g}$, (c) follows from $\nabla_{v}R(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}}))=0$ and the $\mu$-strong convexity of $R$, (d) follows from $c_{n}\geq C_{c}\geq L_{g}\geq\mu$ and $1-x\leq e^{-x}$ for $10<x<$ 1, and (e) follows from (G.2.1).
@@
Denote $\kappa_{g}=\frac{L_{g}}{\mu}$. From [7, Equation (6.19)], given $x_{t}\in{\mathcal{M}}_{x}$ and $y_{t}^{N_{t}}\in{\mathcal{M}}_{y}$, we have
@@
where $\epsilon_{y,v}$ is defined in Lemma G.1. Summing it from $t=0$ to $T-1$, we have
@@
Summing (35) and (36) from 00 to $T-1$, we have
@@
where the last inequality follows from Assumption 3.4 and $a_{t_{1}}\leq C_{a}$, and $F_{0}$ is defined in Lemma G.2. This result is the same as (G.3).
@@ H.1.1 Improvement on objective function for one step update
[…]
@@
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq(1+\frac{1}{b_{k+1}^{2}}\bar{\zeta}L_{g}^{2}-\frac{\mu}{b_{k+1}})d^{2}(y_{t}^{k},y^{*}(x_{t})),$
@@
the third inequality follows from Assumption 3.5 by employing [7, Proposition 10.22], i.e.,
@@
$C_{b}:=\max\{2\bar{\zeta}L_{g}\frac{L_{g}}{\mu},b_{0}\},~{}~{}C_{c}:=\max\{L_{g},c_{0}\},$
@@
If $k_{1}$ in Proposition H.1 does not exist, we have $b_{K_{t}}\leq C_{b}$. By [80, Lemma 2], it holds that $K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $K_{t}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, by $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{x_{t}}^{2}\geq\epsilon_{y}$ and $b_{k}\leq C_{b}$ holds for all $k<K^{t}$, we have
@@
where the first inequality follows from $b_{k}\geq C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$ and $1-m\leq e^{-m}$ for $10<m<$ 1, specifically, we have $C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$ and it holds that $10<\frac{\mu}{b_{k}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{k}^{2}}<$ 1, and the second inequality follows from (H.1.1).
@@ I.1.1 Simple problem
[…]
@@
I.1.2 Deep hyper-representation for classification
@@ I.2.1 Shallow hyper-representation for regression
Figure 4: Shallow hyper-representation for regression (Left: $n=200$, Right: $n=1000$).
@@ I.2.2 Deep hyper-representation for classification
Figure 5: Deep hyper-representation for classification (Left: $n=200$, Right: $n=1000$).
@@
Appendix J Broader impacts and technical limitations
This paper proposes an algorithm for solving Riemannian bilevel optimization problems. No negative societal impacts are anticipated from this research that warrant disclosure.
Notably, this work focuses exclusively on developing an adaptive double-loop algorithm for deterministic Riemannian bilevel optimization (RBO) problems characterized by geodesically strongly convex lower-level objectives, a limitation stemming from the inherent structural complexities of RBO. Future research directions include: (1) designing single-loop adaptive algorithms; (2) extending the framework to stochastic settings via adaptive step sizes using inverse cumulative stochastic hypergradient norms or other strategies [18, 42, 74, 80, 52, 57]; (3) addressing geodesically convex (non-strongly convex) lower-level objectives through regularization [2] or Riemannian Polyak-Łojasiewicz (PŁ) conditions [11]; and (4) investigating alternative adaptive step size strategies to resolve the factor-$1/\epsilon$ gap in gradient ($G_{g}$) and Hessian-vector product ($HV_{g}$) complexities compared to existing Riemannian bilevel methods [29, 48].

### v3

@@ G.2.1 Improvement on objective function for one step update
where (a) follows from Baillon-Haddad Theorem [4, Theorem 5.8 (iv)], i.e., for $y\in{\mathcal{M}}_{y}$ and $v\in T_{y}{\mathcal{M}}_{y}$, it holds that
@@
where (a) follows from (33), (b) follows from $c_{n}>C_{c}\geq L_{g}$, (c) follows from $\nabla_{v}R(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}}))=0$ and the $\mu$-strong convexity of $R$, (d) follows from $c_{n}\geq C_{c}\geq L_{g}\geq\mu$ and $1-x\leq e^{-x}$ for $0<x<1$, and (e) follows from (G.2.1).
@@
Denote $\kappa_{g}=\frac{L_{g}}{\mu}$. From [10, Equation (6.19)], given $x_{t}\in{\mathcal{M}}_{x}$ and $y_{t}^{N_{t}}\in{\mathcal{M}}_{y}$, we have
@@
where $\epsilon_{y,v}$ is defined in Lemma G.1. Summing it from $t=0$ to T1T-1, we have
@@
Summing (36) and (37) from 0 to T1T-1, we have
@@
where the last inequality follows from Assumption 3.4 and $a_{t_{1}}\leq C_{a}$, and $F_{0}$ is defined in Lemma G.2. This result is equivalent to (G.3).
@@ H.1.1 Improvement on objective function for one step update
$\left\{\begin{aligned} a_{t}\leq&C_{a},~~&t\leq t_{1};\\ a_{t}\leq&C_{a}+2F_{0}+\frac{2t\epsilon_{y,v}}{a_{0}},~~&t\geq t_{1},\end{aligned}\right$.
@@
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq\left(1+\frac{1}{b_{k+1}^{2}}\bar{\zeta}L_{g}^{2}-\frac{\mu}{b_{k+1}}\right)d^{2}(y_{t}^{k},y^{*}(x_{t})),$
@@
the third inequality follows from Assumption 3.5 by employing [10, Proposition 10.22], i.e.,
@@
$C_{b}:=\max\left\{2\bar{\zeta}L_{g}\frac{L_{g}}{\mu},b_{0}\right\},~~C_{c}:=\max\{L_{g},c_{0}\},$
@@
If $k_{1}$ in Proposition H.1 does not exist, we have $b_{K_{t}}\leq C_{b}$. By [91, Lemma 2], it holds that $K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $K_{t}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, by $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{x_{t}}^{2}\geq\epsilon_{y}$ and $b_{k}\leq C_{b}$ holds for all $k<K^{t}$, we have
@@
where the first inequality follows from $b_{k}\geq C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$ and $1-m\leq e^{-m}$ for $0<m<1$, specifically, we have $C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$ and it holds that $0<\frac{\mu}{b_{k}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{k}^{2}}<1$, and the second inequality follows from (H.1.1).
@@ I.1.1 Simple problem
For solving the linear system: we adopt the conjugate gradient (CG) method as the default approach in RHGD. For our algorithm, the CG procedure is terminated when the residual norm falls below a tolerance of $10^{-10}$ or the number of CG iterations reaches $50$. Note that the tolerance $10^{-10}$ for CG procedure can be seen as $\mathcal{O}(\frac{1}{T^{\alpha}})$ with $\alpha>1$ is a constant, this will not influence the total iterations and the complexity results since the convergence rate of the CG procedure is linear. Moreover, for the GD procedure, we employ the stopping criteria specified in Algorithm 3 or terminate when the number of iterations reaches $\min\{50\lfloor t/5\rfloor,500\}$, where $\lfloor\cdot\rfloor$ denotes the floor function, and $t$ represents the iteration round $t$. This warm-start-inspired strategy aims to rapidly obtain a reasonable solution during early iterations, and progressively increase internal iterations to satisfy the gradient norm tolerance $\epsilon_{v}$ in later stages.
@@
I.1.2 Robustness analysis
@@ I.2.1 Shallow hyper-representation for regression
Figure 4: Shallow hyper-representation for regression (Left two column: $n=200$, Right two column: $n=1000$).
@@ I.2.2 Deep hyper-representation for classification
Figure 5: Deep hyper-representation for classification (Left: sampling ratio 12.5%, Right: sampling ratio 25%).
@@
I.2.3 Robust optimization on manifolds
In this subsection, following [53, Section 7.1], we consider the robust optimization on manifolds, which have the following forms:
$\begin{array}[]{lcl}&\min\limits_{p\in\Delta_{n}}&\left\|p-\frac{\mathbf{1}}{n}\right\|^{2}-\sum_{i=1}^{n}p_{i}\ell(y;\xi_{i})\\ &{\rm s.t.}&y\in\operatorname*{arg\,min}\limits_{y\in\mathbb{S}_{++}^{d}}\sum_{i=1}^{n}p_{i}\ell(y;\xi_{i}),\end{array}$ (50)
where $\Delta_{n}:=\{p\in\mathbb{R}^{n}:\sum_{i=1}^{n}p_{i}=1,p_{i}>0\}$ and $\mathbb{S}_{++}^{d}:=\{y\in\mathbb{R}^{d\times d}:y\succ{\mathbf{0}}\}$ represent the multinomial manifold (or probability simplex) and positive definite matrix space, respectively. For robust Karcher mean (KM) problem, $\ell(y;\xi_{i})$ in Problem (50) represents the geodesic distance of two positive definite matrices [7]:
$\ell(y;\xi_{i})=\|\log(y^{-1/2}\xi_{i}y^{-1/2})\|^{2}_{F},$
where $\xi_{i}$’s are the symmetric positive definite data matrices. For robust maximum likelihood estimation (MLE) problem, $\ell(y;\xi_{i})$ in Problem (50) represents the log likelihood of the Gaussian distribution [74]:
$\ell(y;\xi_{i})=\frac{1}{2}\log\det(y)+\frac{{\xi_{i}}^{\top}y^{-1}\xi_{i}}{2},$
where $\xi_{i}$’s are sample vectors from the Gaussian distribution.
In this experiment, we set the step sizes as $\eta_{x}=\eta_{y}=1/\alpha_{0}=1/\beta_{0}=1/\gamma_{0}=0.1$, with all other settings identical to those in Section 4.2. For the robust Karcher mean problem, we set $d=20$ and $n\in\{5,10,20,50\}$, randomly generating the symmetric positive definite data matrices $\xi_{i}$’s. For the robust maximum likelihood estimation problem, we set $d=50$ and $n\in\{100,300,500,1000\}$, and randomly generate the sample vectors from a Gaussian distribution with mean $\mathbf{0}$ and covariance matrix being a random positive semidefinite symmetric matrix. We use five different random seeds to conduct experiments for each group. The results are shown in Figure 6, where the x-axis represents the average time, and the y-axis represents the average value of the ergodic performance $\min_{i\in[0,t]}\|\widehat{{\mathcal{G}}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}$. Note that for the MLE experiment with $n=100$, we do not report the results of RHGD-20 because it fails to converge in this case. As shown in Figure 6, compared to RHGD, AdaRHD achieves superior results more efficiently and robustly, which further demonstrates the superiority and robustness of our algorithm.
(a) $n=5$
(b) $n=10$
(c) $n=20$
(d) $n=50$
(e) $n=100$
(f) $n=300$
(g) $n=500$
(h) $n=1000$
Figure 6: Robust optimization on manifolds (Top row: KM model with $d=20$ and different $n$, Bottom row: MLE model with $d=50$ and different $n$).